I thought I was doing pretty good with absolute value, but this one has me stumped. Could you please help me solve this problem?
15-|2+3x| /2=7
thank you in advance for your help

Answer by benni1013(165)

15-|2+3x| /2=7==>Multiply each side by two==>15-|2+3x|=14==>Subtract 15==>
-|2+3x|=-1==>Treat this like multiply by -1 so divide by -1==>|2+3x|=1==>Now separate into a negative case and positive case==>-2-3x=1 and 2+3x=1==>Solve for each==>-2-3x=1==>Add 2==>-3x=3==>x=1==>2+3x=1==>Subtract 2==>3x=-1==>Divide by 3==>x=-1/3